OpenStudy (anonymous):

is the answer (1, -1)? Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1

OpenStudy (anonymous):

@JuanitaM

OpenStudy (anonymous):

@RoseDryer can you help?

OpenStudy (anonymous):

how did you get (1,-1)? i got a different answer, can you show me your solution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you show it to me then?

OpenStudy (anonymous):

give me a second i wrote it in a paper trying to find it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2x + y = 3 -3x + 4y = 1 -x + 5y = 4 and replace 1 by x and get - 1 + 5y = 4 +1 +1 5y = 5 /5 /5 y = 1 and x = 1 <<<< oh is that the answer?

OpenStudy (anonymous):

hum.... i don't think the problem is asking about that Substitution 2x + y =3 and -3x+4y = 1 , if you move 2x from the first eq. you get y = 3 - 2x now you can substitute y in the second eq with y = 3-2x

OpenStudy (anonymous):

can i show you the answer choices and you tel me what you think it is?

OpenStudy (anonymous):

where did you get -x + 5y = 4???

OpenStudy (anonymous):

i think shine added the two eqs

OpenStudy (anonymous):

i put them together the two equations together

OpenStudy (anonymous):

are there three equations?

OpenStudy (anonymous):

aahh.. ok..

OpenStudy (anonymous):

follow @i3lue 's instructions, that's the right steps for substitution..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

answer choices VV (1, 1) (-1, -1) (1, -1) (-1, 1)

OpenStudy (anonymous):

i got (1,1)

OpenStudy (anonymous):

thank you but how? i got (1, 1) too so i guess my way is right

OpenStudy (anonymous):

let me show you